I need to find all real solutions to this equation. Please help.
(y - 2/3)^2 = 4/9

Answer by stanbon(75887)

I need to find all real solutions to this equation. Please help.
(y - 2/3)^2 = 4/9
---
Take the sqrt of both sides to get:
y-(2/3) = +-(2/3)
---
y = 0 or y = 4/3
----------------------
Cheers,
Stan H.
=====================